# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ V Ethical Challenges and Best Practices for LLM Integration in UAVs
This section reviews representative ethical challenges in LAUS and surveys emerging best practices for secure, transparent, and accountable deployment in real-world UAV scenarios.
V-A Securing UAV Operations: Privacy and Data Protection with LLMs
@@
While LLMs enhance autonomy and reasoning in UAV networks, they also introduce security and privacy vulnerabilities, including prompt manipulation, leakage of sensitive sensor data, unauthorized access, and unsafe decision-making under adversarial conditions. When LLMs influence mission planning, resource scheduling, or control actions, such vulnerabilities can result in severe operational failures. LLM security encompasses practices that protect models, data, and outputs across development, deployment, and real-time operation. Core objectives include preventing unauthorized access, mitigating adversarial attacks such as jailbreaking and data poisoning, and ensuring compliance with safety and operational constraints.
@@
Jailbreaking is a particularly critical threat in LAUS. These attacks manipulate prompts to bypass safety constraints and induce unauthorized behavior. White-box attacks assume access to internal model information, enabling gradient- or logit-based manipulation, while black-box attacks rely solely on crafted prompts and are therefore more practical in real-world deployments. Among black-box methods, template completion attacks are especially dangerous. By embedding harmful intent within structured templates, fictional scenarios, or role-playing contexts, attackers exploit LLMs’ pattern-completion behavior while evading keyword-based filters. Context-based manipulation and code injection further amplify these risks, particularly for models relying heavily on contextual coherence and demonstration-based reasoning. These risks are amplified in UAV networks using ICL for real-time decision-making. For example, malicious demonstrations injected into scheduling prompts can bias UAVs toward suboptimal sensing actions, such as prioritizing sensors with poor channel conditions, leading to increased packet loss, degraded network performance, and compromised mission objectives [30]. This illustrates how prompt-level attacks can propagate into system-level failures.
Mitigation requires a multi-layered defense strategy. Integrating an independent safety verifier allows LLM-generated actions to be evaluated against predefined constraints and overridden if unsafe. Complementary attack detection modules can monitor prompt statistics, such as perplexity or structural anomalies, to identify manipulations in real time. Together, these mechanisms enable robust operation under adversarial conditions.
Table XIII: Framework for LLM observability and safety verification in UAV decision-making systems, detailing monitoring mechanisms, traceability, feedback integration, and rule-based verification to ensure transparency, reliability, and operational safety.
Concept Goal Key Mechanism Benefits Notes
LLM Observability Ensure trustworthy, explainable, and consistent outputs Continuous monitoring of model outputs using defined evaluation metrics (correctness, relevance, coherence, factual consistency) Detects model drift, data shifts, hallucinations; enables optimization and transparency Requires systematic logging, monitoring infrastructure; basic metrics alone are insufficient
Prompts and User Feedback Track quality and relevance of LLM outputs Structured logging of prompts, API calls, and user feedback annotated with metadata Identifies prompt failures; optimizes prompt templates; improves user experience Dependence on accurate user feedback; may require additional preprocessing
Tracing Understanding system performance and workflow Captures traces of user interactions, broken down into spans representing workflow steps Pinpoints bottlenecks; shows how components interact; supports debugging Requires proper instrumentation; may add computational overhead
Latency and Usage Monitoring Optimize performance and resource utilization Measure response times, token usage, throughput, and resource consumption Identifies bottlenecks; helps scale efficiently; reduces unnecessary latency Does not capture quality or correctness of outputs by itself
Safety Verification Layer Ensure reliable and safe decision-making Independent verifier applies rule-based scoring against predefined safety constraints; overrides unsafe LLM outputs Guarantees safety compliance; prevents execution of unsafe or illogical actions Introduces additional processing; relies on well-defined safety rules; deterministic rules may limit flexibility
LLM + Safety Verifier Integration Combine reasoning flexibility with safety LLM generates adaptive decisions; safety verifier enforces constraints before execution Balances adaptability and reliability; ensures safe operation in dynamic systems Coordination between LLM and verifier needed; probabilistic LLM outputs still require careful monitoring
@@ V-B Bias-Aware LLMs for Safer UAV Operations and Communications
Bias in LLMs arises from imbalanced or unrepresentative training data and can directly affect safety and mission effectiveness in UAV systems, for example through skewed threat assessment or unfair scheduling decisions. These risks are particularly severe in post-disaster scenarios characterized by uncertainty, limited communication, and strict latency constraints. In large-scale SAR missions, relying on a single monolithic LLM for perception, planning, and scheduling can create an operational bottleneck. Under high cognitive load, such models are more prone to hallucinations, biased prioritization, and delayed responses, violating key requirements for multi-UAV operations, including real-time responsiveness, efficient resource use, and robustness to partial failures.
@@
Beyond architecture, LLM observability plays a critical role in bias mitigation by enabling inspection of intermediate reasoning steps and decision patterns in real time [35]. While observability does not eliminate bias at its source, it provides a practical mechanism for detecting and correcting biased behavior in deployed systems.
Complementing observability, Morales et al.[86] propose LangBiTe, a model-driven testing framework that enables continuous ethical assessment through domain-specific specifications and automated scenario generation. Such structured testing provides additional assurance for bias detection and accountability in LAUS.
@@ V-C Ethical LLMs in Action: Ensuring Accountability and Transparency for UAV Operation and Communication
This principle is exemplified by frameworks such as FlyAdapt, which integrates an LLM-assisted causal reasoning engine with predictive edge caching and hybrid verification for multi-UAV systems. By prioritizing interpretability and real-time validation, FlyAdapt achieves substantial latency reduction while intercepting the majority of diverse attacks, demonstrating how transparent, modular design can support reliable deployment in high-stakes environments. To manage residual risks, LLM observability must be adopted as a foundational practice. Observability connects input quality to output accountability through continuous monitoring, automated evaluation, and human feedback across the AI lifecycle. These mechanisms mitigate hallucination and performance drift, transforming LLMs from opaque components into governable systems.
@@ V-D Sustainable Intelligence: Environmental Concerns in LLMs
From an energy perspective, training dominates carbon emissions, with large-scale models consuming electricity comparable to hundreds of households, while inference incurs smaller per-query costs. However, at the scale of billions of daily interactions, inference energy consumption becomes a significant contributor to global emissions. Environmental impacts are further amplified by lifecycle factors such as resource-intensive hardware manufacturing, embodied carbon in AI accelerators, and regional variation in data-center energy mixes. Design trade-offs also matter: larger models may be more efficient per high-quality output, whereas smaller models can be preferable under high request volumes.
@@
V-E Key Findings and Insights
Recent studies consistently show that effective integration of LLMs into UAV systems depends not only on model capabilities but also on system-level mechanisms that ensure reliability, safety, and accountability. In practice, this calls for a layered governance framework combining LLM observability, explicit safety verification, and HITL oversight.
LLM observability forms the foundation of trustworthy deployment by linking output quality with accountability. Unlike traditional monitoring focused on latency or resource usage, observability enables continuous assessment of LLM outputs across correctness, relevance, coherence, and factual consistency, allowing early detection of model drift, hallucinations, and performance degradation in dynamic UAV environments [12]. Modern observability extends classical logs and traces with LLM-specific signals, including structured prompt logging, user feedback, request tracing across multi-stage pipelines, and token-usage monitoring, supporting systematic diagnosis and scalable optimization [89], [88].
However, observability alone is insufficient to guarantee safety. Due to their probabilistic nature, LLMs may still produce decisions that violate operational or safety constraints. A second key insight is therefore the necessity of an independent safety verification layer, where LLM-generated decisions are evaluated against predefined rules and feasibility constraints and overridden when unsafe. This separation of adaptive reasoning and deterministic enforcement enables UAV systems to combine flexibility with the reliability required for safety-critical operation. Representative observability and verification components are summarized in Table XIII.
Beyond automated safeguards, HITL oversight remains indispensable. While LLMs provide strong reasoning and decision-support capabilities, they are inherently limited in handling ethical dilemmas, rare edge cases, and ambiguous real-world scenarios. HITL frameworks integrate human judgment into mission planning and real-time operation, allowing operators to intervene when uncertainty or risk is high [108]. Effective HITL systems balance automation with oversight by reserving human intervention for complex or high-stakes decisions, supported by clear role definitions, intuitive interfaces, escalation criteria, and continuous feedback loops. Diversity among human reviewers further reduces bias and strengthens accountability [69], [26].
Taken together, these findings indicate that robust LAUS should be designed as governed socio-technical systems rather than standalone intelligent agents. LLM observability provides transparency and quality control, safety verifiers enforce hard constraints, and HITL mechanisms supply ethical judgment and resilience, enabling UAV systems to scale intelligence while preserving safety, trust, and accountability in real-world deployment.
VI ICL-Based Case Study for Multi-UAV Systems
This section transitions from theoretical frameworks to practical validation by presenting an original case study on ICL for multi-UAV coordination. We demonstrate the application of an edge-hosted LLM to jointly optimize data-collection scheduling and velocity control in a dynamic emergency-response scenario. The case study details the system architecture, simulation environment, and operational workflow of the proposed ICL-based framework. Finally, we present comparative performance results against traditional RL baselines, empirically highlighting the advantages of LLM-assisted adaptability and real-time decision-making in UAV networks.
This case study applies the ICL framework to multiple UAVs that jointly optimize data collection scheduling and velocity control. When multiple UAVs operate simultaneously, the actions of each UAV influence not only its own future states but also the behavior of other UAVs, resulting in a coupled multi-agent decision problem.
In this scenario, each UAV interacts with an unknown environment and periodically transmits collected data and state information to an LLM. The LLM uses four key components – the task description ($TD_{\text{task}}$), example set ($Ex_{t}$), current environment state ($e_{t}$), and its internal reasoning – to generate the next action ($a_{t}$). The task description encodes objectives, constraints, and expected outcomes, enabling operators to avoid explicitly designing optimization models. During operation, the LLM iteratively updates its decisions based on environment feedback, example demonstrations, and prior UAV actions, gradually converging on an effective policy.
VI-A Multi-UAV Interaction and Decision Coupling
(a) DCS-ICL vs MADQN packet loss comparison
(b) Performance comparision of different LLMs.
Figure 10: Comparative analysis of packet loss over time steps between the proposed DCS-ICL method and the baseline MADQN approach along with performance evaluation of different LLMs.
In the multi-UAV scenario, each UAV independently selects sensors and adjusts its velocity. However, these decisions are interdependent and require a multi-agent ICL formulation to jointly minimize the total system cost, accounting for factors such as packet loss, buffer overflow, and failed transmissions.
We propose the ICL-based Data Collection Scheduling and Velocity Control (DCS-ICL) framework for UASNets. DCS-ICL leverages ICL to optimize emergency data collection schedules and UAV velocities. Specifically, each UAV gathers sensory information, including queue lengths, battery levels, channel conditions, and UAV locations, from ground sensors and transmits this data to an edge-hosted LLM. The LLM is initialized with a structured system prompt that defines its operational role, constraints, and ethical safeguards, thereby ensuring compliance with safety and security requirements.
Upon receiving environmental data, the LLM generates a detailed task description in natural language that includes mission objectives, input parameters, operational rules, illustrative examples, expected outputs, and a feedback mechanism. In this context, the LLM infers an optimal data-collection schedule and velocity, which the UAVs then execute. Performance metrics and updated environmental states are continuously monitored and fed back into subsequent decision cycles, enabling dynamic adaptation and improvement of scheduling effectiveness over time.
VI-B Simulation Setup
We consider a scenario in which 10 ground sensors are randomly placed within a $100\penalty 10000\ \text{m}\times 100\penalty 10000\ \text{m}$ area. Each sensor is configured with a maximum battery capacity of $50\penalty 10000\ \text{J}$, a transmit queue limit of 40 packets, and a maximum transmit power of $100\penalty 10000\ \text{mW}$. In the learning configuration, each episode consists of 30 time steps. Simulations are performed in a customized Python 3.5 environment. UAV behavior is modeled according to the policies learned by the DCS-ICL framework. The wireless connection between UAVs and sensors is implemented using the probabilistic channel model. All experiments are conducted on a Lenovo workstation running Ubuntu 20.04 LTS, equipped with an Intel Core i5-7200U CPU @ 2.50 GHz × 4 and 16 GB of RAM. The baseline technique is Multi-agent Deep-Q-Network (MADQN): Each UAV optimizes its data collection schedule and velocity using DQN by dynamically adapting to key environmental conditions, including UAV locations, queue length, battery level, and channel conditions, to minimize packet loss.
VI-C Performance Evaluation
Fig. 10(a) illustrates the convergence of the proposed approach using GPT-4O-mini, compared to the MADQN baseline (MADQN performance is reported for the last episode). Both the proposed approach and MADQN achieve comparable performance. Fig. 10(b) illustrates the temporal evolution of packet loss for the compared models. As observed, all methods start with a relatively high packet loss at the initial time step. However, Grok and the two Mistral-based models exhibit rapid convergence, reducing packet loss to zero within the first two to three time steps and maintaining stable zero-loss performance thereafter. In contrast, LLMA shows only a partial reduction, stabilizing at a non-zero packet loss level of approximately eight packets and failing to achieve full convergence throughout the evaluation horizon. These results demonstrate the superior convergence speed and stability of the proposed approach relative to LLMA, highlighting its effectiveness in quickly mitigating packet loss and sustaining reliable network performance.
VII Most recent trends
The integration of LLMs into specialized domains, such as UAV networks, presents both unprecedented opportunities and complex challenges. Although current models demonstrate remarkable capabilities in natural language understanding, reasoning, and autonomous decision-making, limitations persist in areas such as domain-specific expertise, efficiency, interpretability, and sustainability. Addressing these gaps requires targeted research that improves model performance while ensuring practical, ethical, and environmentally conscious deployment. The following directions highlight key avenues for advancing LLMs in UAVs and related applications, including improvements in specialization, efficiency, privacy, reliability, and integration with emerging technologies such as AGI and quantum computing.
VII-A Near-Term Directions
VII-A1 Domain-Specific UAV LLMs for Autonomous Aerial Intelligence
A salient direction for future research is the development of fine-tuned, domain-specific LLMs explicitly designed for UAV applications. Drawing on industry trends in finance and healthcare, a specialized ”UAV-LLM” would be pretrained on aeronautical manuals, flight regulations, and telemetry data, then fine-tuned for mission-specific tasks such as trajectory planning, anomaly detection, and real-time decision-making. This deep specialization is expected to reduce hallucinations, improve accuracy, and provide a reliable cognitive core for UAV autonomy, enhancing operational safety, mission efficiency, and natural language interaction with ground control.
Future research should also address the environmental and economic costs associated with large-scale LLMs. Rather than indiscriminate parameter scaling, efforts should focus on domain-specific optimization, carbon accountability, and strategic investment in practical AI applications. By combining specialized model development with sustainable and cost-effective practices, UAV LLMs can deliver high performance and responsible deployment, ensuring scalable, trustworthy, and environmentally conscious aerial intelligence systems.
VII-A2 Small and Edge-Deployable LLMs for UAV Applications
@@
VII-B Medium-Term Directions
VII-B1 Self-Improving LLMs for UAV Applications
A promising direction for future research is the development of self-improving LLMs capable of generating synthetic training data to address the critical shortage of domain-specific datasets in the UAV sector. By autonomously creating and fine-tuning on curated, self-generated question-answer pairs, these models can achieve significant performance gains. In the UAV context, this approach could enable the generation of large-scale datasets covering operational scenarios, including navigation commands, anomaly detection procedures, and sensor data interpretation. Leveraging self-training capabilities enables continuous refinement of domain expertise, reduces reliance on costly manual data collection, and accelerates the development of highly specialized, efficient LLMs. Research in this area could unlock more advanced, autonomous aerial capabilities and serve as a blueprint for applying self-improving models in other data-scarce domains.
VII-B2 Reducing Hallucinations in LLMs
@@ VII-B3 Towards Reliable LLM Systems:
Research in these areas will shift the focus from building ”bigger” models to creating trustworthy, production-ready LLM systems, enabling broader adoption in UAV networks while mitigating inherent limitations. The ultimate goal is to develop LLM-assisted systems that are not only fluent and capable but also reliable, safe, and self-regulating within a larger operational context.
VII-C Long-term Directions
VII-C1 AGI-Enhanced LLMs
Advancing LLMs with AGI-inspired enhancements offers a promising direction for future research. Although transformer-based LLMs have demonstrated impressive abilities in language understanding and generation, they remain limited in reasoning, adaptability, and real-world comprehension. Research aimed at integrating AGI-driven cognitive frameworks could enable LLMs to incorporate symbolic reasoning, causal inference, and dynamic knowledge updates, supporting deeper conceptual understanding beyond statistical pattern recognition. These advancements could also improve bias mitigation, ethical governance, and multimodal reasoning, ultimately allowing LLMs to generalize knowledge more effectively across the UAV domain.
VII-C2 Sustainable LLM Infrastructure
The high computational and energy demands of training and deploying LLMs remain a major constraint in the global AI landscape. For example, LLaMA-3-70B emits approximately 13 kg of $CO_{2}$ per 1 million-token batch, while GPT-3’s training consumed 1,287 MWh – enough to power 120 U.S. households for a year. Inference adds further energy load, accounting for roughly 60% of total ML energy consumption. Regional factors, such as water-intensive cooling in Arizona or fossil-fuel-heavy electricity grids in Chile, amplify environmental impacts. By 2030, AI could consume up to 10% of global electricity, exceeding the entire output of some countries, including Belgium.
VIII Conclusion
The integration of LLMs and MLLMs into UAV systems marks a significant step toward intelligent and adaptive aerial operations. This survey examined the convergence of LLM technologies with UAV communications, control, and ethical frameworks, providing a unified view of system architectures, methodologies, and real-world applications. By enabling advanced reasoning, natural language interaction, and context-aware decision-making, LLMs support applications ranging from real-time SAR to swarm coordination and mission planning. Techniques such as RAG and prompt engineering allow UAVs to operate dynamically in complex environments without extensive retraining, while MLLMs further extend these capabilities through multimodal perception and human-swarm interaction. Despite their potential, LLM-assisted UAV systems raise critical challenges related to bias, transparency, accountability, and environmental impact. Addressing these issues requires robust governance mechanisms, including HITL oversight and LLM observability, to ensure safe and trustworthy deployment. Looking forward, key research directions include domain-specific UAV LLMs, efficient edge deployment, hallucination reduction, enhanced multimodal reasoning, and sustainable AI practices. The continued synergy among LLMs, 5G/6G connectivity, edge AI, and emerging computing paradigms is expected to accelerate the evolution of intelligent UAV ecosystems, enabling more reliable and responsible autonomous aerial systems for societal applications.

### v2

@@ V Ethical Challenges and Future Directions
This section reviews representative ethical challenges and future directions for secure, transparent, and accountable deployment in real-world UAV scenarios.
V-A Securing UAV Operations: Privacy and Data Protection
@@
While LLMs enhance autonomy and reasoning in UAV networks, they also introduce security and privacy vulnerabilities, including prompt manipulation, leakage of sensitive sensor data, unauthorized access, and unsafe decision-making under adversarial conditions. When LLMs influence mission planning, resource scheduling, or control actions, these vulnerabilities can lead to severe operational failures. LLM security encompasses practices that protect models, data, and outputs across development, deployment, and real-time operation. Core objectives include preventing unauthorized access, mitigating adversarial attacks such as jailbreaking and data poisoning, and ensuring compliance with safety and operational constraints.
@@
Jailbreaking is a particularly critical threat in LAUS. These attacks manipulate prompts to bypass safety constraints and induce unauthorized behavior. White-box attacks assume access to internal model information, enabling gradient- or logit-based manipulation, while black-box attacks rely solely on crafted prompts and are therefore more practical in real-world deployments. Among black-box methods, template-completion attacks are particularly dangerous. By embedding harmful intent within structured templates, fictional scenarios, or role-playing contexts, attackers exploit LLMs’ pattern-completion behavior while evading keyword-based filters. Context-based manipulation and code injection further amplify these risks, particularly for models relying heavily on contextual coherence and demonstration-based reasoning. These risks are amplified in UAV networks using ICL for real-time decision-making. For example, malicious demonstrations injected into scheduling prompts can bias UAVs toward suboptimal sensing actions, such as prioritizing sensors with poor channel conditions, thereby increasing packet loss, degrading network performance, and compromising mission objectives [29]. This illustrates how prompt-level attacks can propagate into system-level failures.
@@ V-B Bias-Aware LLMs for Safer UAV Operations and Communications
Bias in LLMs arises from imbalanced or unrepresentative training data and can directly affect safety and mission effectiveness in UAV systems, for example, through skewed threat assessment or unfair scheduling decisions. These risks are particularly severe in post-disaster scenarios characterized by uncertainty, limited communication, and strict latency constraints. In large-scale SAR missions, relying on a single monolithic LLM for perception, planning, and scheduling can create an operational bottleneck. Under high cognitive load, such models are more prone to hallucinations, biased prioritization, and delayed responses, violating key requirements for multi-UAV operations, including real-time responsiveness, efficient resource use, and robustness to partial failures.
@@ V-C Ethical LLMs in Action: Ensuring Accountability and Transparency for UAV Operation and Communication
This principle is exemplified by frameworks such as FlyAdapt, which integrates an LLM-assisted causal reasoning engine with predictive edge caching and hybrid verification for multi-UAV systems. By prioritizing interpretability and real-time validation, FlyAdapt achieves substantial latency reduction while intercepting the majority of diverse attacks, demonstrating how a transparent, modular design can support reliable deployment in high-stakes environments. To manage residual risks, LLM observability must be adopted as a foundational practice. Observability connects input quality to output accountability through continuous monitoring, automated evaluation, and human feedback across the AI lifecycle. These mechanisms mitigate hallucination and performance drift, transforming LLMs from opaque components into governable systems.
@@ V-D Sustainable Intelligence: Environmental Concerns in LLMs
From an energy perspective, training dominates carbon emissions, with large-scale models consuming electricity comparable to that of hundreds of households, whereas inference incurs lower per-query costs. However, at the scale of billions of daily interactions, inference energy consumption becomes a significant contributor to global emissions. Environmental impacts are further amplified by lifecycle factors, including resource-intensive hardware manufacturing, embodied carbon in AI accelerators, and regional variation in data-center energy mixes. Design trade-offs also matter: larger models may be more efficient per high-quality output, whereas smaller models can be preferable under high request volumes.
@@
V-E Small and Edge-Deployable LLMs for UAV Applications
@@
V-F Reducing Hallucinations Towards Reliable LLM Systems
@@
VI Conclusion
The integration of LLMs and MLLMs into UAV systems represents a significant step toward intelligent, adaptive aerial operations. This survey examined the convergence of LLM technologies with UAV communications, control, and ethical frameworks, providing a unified view of system architectures, methodologies, and real-world applications. By enabling advanced reasoning, natural language interaction, and context-aware decision-making, LLMs support applications ranging from real-time SAR to swarm coordination and mission planning. Techniques such as RAG and prompt engineering allow UAVs to operate dynamically in complex environments without extensive retraining, while MLLMs further extend these capabilities through multimodal perception and human-swarm interaction. Despite their potential, LLM-assisted UAV systems pose critical challenges regarding bias, transparency, accountability, and environmental impact. This work is expected to accelerate the evolution of intelligent UAV ecosystems, enabling more reliable and responsible autonomous aerial systems for societal applications.